OpenStudy (anonymous):

Simplify (a^8a^6)^1/7 /a^2

OpenStudy (geekfromthefutur):

Hey @Kdgriff2 welcome to openstudy!

OpenStudy (anonymous):

-Thanks! :)

OpenStudy (geekfromthefutur):

Welcome!

OpenStudy (geekfromthefutur):

Ok so lets see here

OpenStudy (geekfromthefutur):

Grab your caculater if you dont have one grab one from your star menu on ur computer

OpenStudy (anonymous):

Ok

OpenStudy (geekfromthefutur):

Sorry its taken so long i am tryin to fgure out how to get it workin

OpenStudy (anonymous):

Its fine

OpenStudy (geekfromthefutur):

There's your problem right there

OpenStudy (anonymous):

Thank you so much!

OpenStudy (geekfromthefutur):

So what you need to do with your calculater is add 8 + 6 and tell me what you get

OpenStudy (geekfromthefutur):

8+6= __ if you have to do it with your fingers.

OpenStudy (geekfromthefutur):

Thats not the answer lol thats your equation in writtin form just answer this and you have your answer

TheSmartOne (thesmartone):

\[\frac{\sqrt[7]{a^8a^6} }{ a^2 }=\frac{ \sqrt[7]{a ^{8+6}} }{ a^2 }=\frac{ \sqrt[7]{a ^{14}} }{ a^2 }=\frac{ a^{{14}\times \frac{ 1 }{ 7 } }}{ a^2 }=\frac{ a^2 }{ a^2 }=1\]